# EnScale: Temporally-consistent multivariate generative downscaling via proper scoring rules

### v2

@@ 7 Discussion
In this study, we present EnScale, a method for emulating RCMs. We compare EnScale to various benchmarks, both physically-motivated (analogues) and existing state-of-the-art machine learning methods (GAN and a diffusion model, CorrDiff). We find that EnScale achieves great performance in terms of a wide range of metrics, both when evaluated per grid point and per variable separately, as well as for spatial structure and dependencies between variables. To our knowledge, EnScale-t is the first downscaling approach that models temporal dependencies in the high-resolution output, improving the temporal consistency of the output substantially compared to all other benchmarks.
@@
Compared to physics-based climate models as well as state-of-the-art machine learning models, our emulator is computationally very efficient. Training EnScale completes in approximately one day. Afterwards, generating a downscaled ensemble with ten members for 100 years of GCM data is completed in under two hours. CorrDiff instead needs about 10 days for training and 36 hours for inference. Running the RCM for the same period would typically take weeks and still yield only one ensemble member. The sparse architecture in EnScale is efficient and scalable also for larger domains and more climate variables.
@@
Our work established a comprehensive evaluation framework of generative downscaling approaches for calibration, spatial and temporal structure, extremes, and multivariate dependencies. Notably, evaluating such methods is not a straightforward task. The output comprises many locations, variables and time steps, and capturing all these aspects is challenging [Maraun2015VALUEStudies]. Model evaluation and statements about skill are necessarily specific for each question and application. So far, our evaluation methods are primarily statistically motivated and inspired by the field of forecast evaluation. There remain open questions, in particular for evaluating the dependencies for many variables at once and for the temporal structure. Evaluating emulators in application-relevant case studies is an interesting avenue of further research [Isphording2024ASimulations].
@@
We found large conditional variability in RCM fields given GCM inputs, arising from internal variability within RCMs. This highlights the value of stochastic emulators: by generating an ensemble of many RCM trajectories for one given driving GCM run, they can capture this internal variability and characterize its spread. This can enable better uncertainty quantification in high-resolution predictions, analogous to the benefits of large ensembles from global climate models [Deser2020InsightsProspects, Rampal2025DownscalingExtremes].
EnScale can readily be applied to data from the three GCMs that were used during training. It can also help to fill the GCM-RCM data matrix, provided that individual GCMs and RCMs were present in the training data. For example, it is possible to sample data from MIROC5-ALADIN63, even though this pair was not present in the training data. We also plan to provide downscaled CMIP6 data after re-training EnScale on more GCMs in CMIP5.
@@
Our downscaling via coarse correction is similar to some existing works: [PriceStephanRasp2022IncreasingModels] also split downscaling of precipitation with a GAN into two steps: a coarse correction step and a pure super-resolution step. The target is radar measurements. Their setup is simpler than ours because they only use data from one numerical model, but still they find that isolating the correction step in a low-dimensional task improves the results. [Wan2024StatisticalModels] downscale data from a single GCM stochastically, considering historical GCM data and using ERA5 reanalysis as a target. Hence, their first step corrects large-scale mismatches between GCMs and ERA5. In a similar fashion, CorrDiff’s mean-residual split (Sec. 4.1.5) corrects some mismatches between input and target data in a simpler task (the mean prediction, first step), before learning the correct variability (residual prediction, second step).
@@
We choose to learn the full path from GCM to RCM, which differs from RCM-emulation in [Doury2024OnPrecipitation, Doury2022]. Previous work argues that emulators which learn a map from GCM to RCM data might pick up patterns specific to each GCM-RCM pair, and are thus less transferable to other GCMs [Hernanz2023OnIssues, Rampal2024EnhancingLearning, Bano-Medina2023TransferabilityApplications]. However, our setup is different because we (i) train with multiple GCM-RCM pairs, possibly enabling the model to learn shared properties across different climate models, and (ii) train a generative model rather than a deterministic one. Future work could also include extrapolation to unseen GCMs.
@@
EnScale leverages both past and future data, enabling it to learn future relationships between low-resolution predictors and high-resolution targets directly from physics-based models. In contrast, model training in perfect prognosis downscaling relies on historical data and aims to extrapolate these learned relationships to future conditions. RCM emulation like in EnScale can readily be combined with a state-of-the art bias-correction method to obtain “future observations”. A promising direction for further development is to explore more sophisticated bias correction algorithms that preserve multivariate structure across time, space, and several variables or integrate observational data also in model training [Hess2025FastLearning, Schmidt2025ASimulation].
@@
EnScale is computationally efficient and can be re-trained also on other datasets from other GCMs and RCMs or for other regions. Also, our emulator could be used to downscale to a higher resolution, for example approximating convection-permitting models as a target. These benefits establish EnScale as a powerful tool for data-driven future climate projections on a local scale and open new opportunities for climate impact studies, where reliable high-resolution projections are essential.
@@ Acknowledgments
Maybritt Schillinger, Maxim Samarin, Nicolai Meinshausen, and Reto Knutti are part of SPEED2ZERO, a Joint Initiative co-financed by the ETH Board. We thank Ruth Lorenz and Urs Beyerle for their help with the data-preprocessing and Sebastian Sippel for valuable inputs during the early stages of this project. Further, we benefitted from many helpful and inspiring discussions with numerous people, among them were Antoine Doury, Jose González-Abad, Shahine Bouabid, Sam Allen, Anna Merrifield, Dominik Schumacher, Daniele Nerini, Niklas Boers and his group. Last but not least, we thank Nina Effenberger and Anya Fries for their constructive feedback on the draft which greatly improved the readability of this work.
Code Availability
@@ Appendix A Additional results
A.1 Distribution shifts between different RCMs
Figure A.1: As Fig. 5, but only comparing to EnScale-no-label and EnScale-t-no-label which exclude the one-hot-encoded index of the GCM and RCM also in the super-resolution model.
We trained two versions of the super-resolution model from coarsened RCM data to the high-resolution target. Once, we included a one-hot-encoded index for the RCM as an additional predictor (as in EnScale, EnScale-t), and once we excluded the index (version EnScale-no-label, EnScale-t-no-label). The index allows the ML model to adjust the output accordingly, dependent on the distribution of this particular RCM. Fig. A.1 summarizes the performance of EnScale-no-label, EnScale-t-no-label compared to EnScale. We find that adding the label only has a very small effect on the results. EnScale often yields small improvements compared to EnScale-no-label, in particular for extremes. This indicates that the step from coarsened RCM data to RCM data is indeed very similar between the different RCMs. For comparison, for the step from GCM to coarsened RCM, including the indices of the RCM improves the energy scores loss by about 8 %.
A.2 Comparison of CorrDiff and CorrDiff-s
CorrDiff (following [Mardani2025ResidualDownscaling] and the implementation in NVIDIA’s PhysicsNeMo library [^3] Code: https://github.com/NVIDIA/physicsnemo/tree/main/examples/weather/corrdiff) generates each ensemble member by drawing a single random noise realization that is shared across all time points. This construction implicitly imposes temporal dependence in the generated time series, but variability across realizations is limited. For this reason, we also tested an adapted version which we call CorrDiff-s. In this variant, we randomize the sampling scheme: For each day, noise is drawn independently of previous days and of other ensemble members. This version represents the variance of samples more faithfully. However, it removes temporal autocorrelation, leading to negative errors in autocorrelations, similar to the other benchmarks (not shown).
CorrDiff-s outperforms CorrDiff in almost all categories (Fig. A.2). As one example, we show that rank histograms by CorrDiff-s are much closer to calibrated (Fig. 8).
We suspect that these discrepancies are a result of large conditional variance in the distribution of RCM fields given GCM inputs. In this case, the original CorrDiff sampling scheme is suboptimal, whereas an independent sampling strategy better captures the spread of possible high-resolution fields.
Figure A.2: As Fig. 5, but comparing CorrDiff, CorrDiff-s, EnScale and EnScale-t.
Figure A.3: As Fig. 8, but comparing CorrDiff, CorrDiff-s, EnScale and EnScale-t.
A.3 Further exemplary samples
@@
Figure A.8: Examples for CorrDiff: Like Fig. 7, but showing examples for the CorrDiff benchmark.
@@
A.4 Results on extrapolation test period
@@
NN-det-extra 4.4 ± 0.08 240.1 ± 1.9 240.0 0.0 1.51 ± 0.01 1.51 ± 0.01 0.8 ± 0.01
EasyUQ-extra 6.3 ± 0.08 179.7 ± 1.5 305.4 251.3 1.20 ± 0.01 2.14 ± 0.02 6.1 ± 0.01
Analogues-extra 8.0 ± 0.13 248.8 ± 2.7 415.5 332.8 1.74 ± 0.01 1.73 ± 0.01 1.2 ± 0.01
GAN-extra 9.3 ± 0.1 223.4 ± 1.42 377.4 308.1 1.48 ± 0.01 1.45 ± 0.01 2.2 ± 0.01
CorrDiff-extra 6.8 ± 0.11 219.5 ± 2.4 393.7 348.3 1.09 ± 0.01 1.08 ± 0.01 0.9 ± 0.01
EnScale-extra 4.4 ± 0.07 166.8 ± 1.57 348.5 363.5 1.18 ± 0.01 1.16 ± 0.01 1.3 ± 0.01
EnScale-t-extra 5.3 ± 0.09 178.5 ± 1.82 362.1 367.2 1.27 ± 0.01 1.26 ± 0.01 1.3 ± 0.01
@@
NN-det-extra 35.1 ± 0.84 631.7 ± 6.64 632.3 0.0 2.38 ± 0.02 6.41 ± 0.07 10.5 ± 0.13
EasyUQ-extra 54.3 ± 0.92 470.0 ± 4.89 838.7 737.2 1.98 ± 0.02 11.39 ± 0.08 9.6 ± 0.05
Analogues-extra 45.7 ± 0.91 567.0 ± 6.95 936.3 738.4 2.43 ± 0.02 6.29 ± 0.06 5.7 ± 0.08
GAN-extra 53.0 ± 1.18 481.5 ± 5.74 798.9 635.5 1.98 ± 0.02 5.39 ± 0.06 8.8 ± 0.08
CorrDiff-extra 55.2 ± 0.96 438.6 ± 4.85 833.7 787.5 1.91 ± 0.02 4.91 ± 0.05 4.0 ± 0.06
EnScale-extra 36.2 ± 0.8 456.4 ± 4.98 844.8 776.5 1.92 ± 0.02 5.05 ± 0.05 5.5 ± 0.07
EnScale-t-extra 36.2 ± 0.81 453.7 ± 5.08 850.1 790.5 1.94 ± 0.02 5.09 ± 0.05 5.3 ± 0.06
@@
NN-det-extra 2.6 ± 0.03 196.4 ± 1.01 196.5 0.0 1.13 ± 0.01 1.44 ± 0.01 2.0 ± 0.02
EasyUQ-extra 4.2 ± 0.03 141.2 ± 0.79 255.6 229.0 0.88 ± 0.0 2.22 ± 0.01 8.9 ± 0.03
Analogues-extra 3.9 ± 0.04 180.8 ± 1.49 304.1 246.4 1.13 ± 0.01 1.39 ± 0.01 2.0 ± 0.02
GAN-extra 4.7 ± 0.04 151.5 ± 0.82 269.9 236.9 0.95 ± 0.0 1.16 ± 0.01 3.6 ± 0.03
CorrDiff-extra 3.8 ± 0.04 137.3 ± 0.84 269.5 264.2 0.82 ± 0.0 1.00 ± 0.01 1.4 ± 0.01
EnScale-extra 2.6 ± 0.03 133.1 ± 0.87 264.4 262.7 0.85 ± 0.0 1.05 ± 0.01 1.8 ± 0.02
EnScale-t-extra 2.7 ± 0.03 134.8 ± 0.91 265.1 260.5 0.86 ± 0.0 1.06 ± 0.01 1.8 ± 0.02

### v3

@@ 7 Discussion
In this study, we present EnScale, a method for emulating RCMs. We compare EnScale to various benchmarks, both physically-motivated (analogues) and machine learning methods (GAN and a diffusion model, CorrDiff(-s)). We find that EnScale achieves competitive performance compared to the state-of-the-art diffusion model w.r.t. a wide range of metrics, both when evaluated per grid point and per variable separately, as well as for spatial structure and dependencies between variables. To our knowledge, EnScale-t is the first downscaling approach to model temporal dependencies in the high-resolution output, approximating temporal consistency of dynamical downscaling. This improves the temporal consistency of the output substantially compared to all other benchmarks.
@@
Compared with physics-based climate models as well as state-of-the-art machine learning models, our emulator is highly computationally efficient. Training EnScale completes in approximately one day. Afterwards, generating a downscaled ensemble with ten members for 100 years of GCM data is completed in under two hours. CorrDiff instead needs about 10 days for training and 36 hours for inference. Running the RCM for the same period would typically take weeks and still yield only one ensemble member. The sparse architecture in EnScale is efficient and scalable also for larger domains and more climate variables.
@@
Our work presented a comprehensive evaluation framework of generative downscaling approaches for calibration, spatial and temporal structure, extremes, and multivariate dependencies. Notably, evaluating such methods is not a straightforward task. The output comprises many locations, variables and time steps, and capturing all these aspects is challenging [41]. Model evaluation and statements about skill are necessarily specific for each question and application. So far, our evaluation methods are primarily statistically motivated and inspired by the field of forecast evaluation. There remain open questions, in particular for evaluating the dependencies for many variables at once and for the temporal structure. Evaluating emulators in application-relevant case studies is an interesting avenue of further research [30].
@@
We found large conditional variability in RCM fields given GCM inputs, arising from internal variability within RCMs. This highlights the value of stochastic emulators: by generating an ensemble of many RCM trajectories for one given driving GCM run, they can capture this internal variability and characterize its spread. This can enable better uncertainty quantification in high-resolution predictions, analogous to the benefits of large ensembles from global climate models [12, 50].
@@
Our downscaling via coarse correction is similar to some existing works: [48] also split downscaling of precipitation with a GAN into two steps: a coarse correction step and a pure super-resolution step. The target is radar measurements. Their setup is simpler than ours because they only use data from one numerical model, but still they find that isolating the correction step in a low-dimensional task improves the results. [74] downscale data from a single GCM stochastically, considering historical GCM data and using ERA5 reanalysis as a target. Hence, their first step corrects large-scale mismatches between GCMs and ERA5. In a similar fashion, CorrDiff’s mean-residual split (Sec. 4.1.5) corrects some mismatches between input and target data in a simpler task (the mean prediction, first step), before learning the correct variability (residual prediction, second step).
We trained our method on a combined dataset, comprising multiple GCMs and RCMs, while conditioning on the RCM in question. This allows the model to benefit from a large dataset and learn shared properties. However, a potential limitation is that the model might regress toward the multi-model mean, underestimating differences among RCMs and failing to capture all model-specific details.
@@
We choose to learn the full path from GCM to RCM, which differs from RCM-emulation in [14, 13]. Previous work argues that emulators which learn a map from GCM to RCM data might pick up patterns specific to each GCM-RCM pair, and are thus less transferable to other GCMs [27, 51, 4]. However, our setup is different because we (i) train with multiple GCM-RCM pairs, possibly enabling the model to learn shared properties across different climate models, and (ii) train a generative model rather than a deterministic one.
EnScale can readily be applied to emulate the three RCMs that were used during training for any GCM input. It can be used to downscale data from CMIP6 for the three GCMs in the train data [15]. As a preliminary result, we also tested extrapolation to two unseen GCMs, HadGEM2-ES and CanESM2, and present the results in Sec. B. For most variables, scores on the unseen GCMs are within the range of scores on the training pairs. Temperature is the only exception, where scores on new GCMs are higher than scores on all GCM-RCM pairs in the training data. We find that, even within the training dataset, scores can differ substantially both between different driving GCMs as well as between different RCMs for the same driving GCM. Hence, studying these differences is an interesting direction of future research to assess the extrapolation scores more rigorously. In summary, however, these results show promising capabilities of EnScale and suggest that the learned map can be generalized across GCMs.
@@
EnScale leverages both past and future data, enabling it to learn future relationships between low-resolution predictors and high-resolution targets directly from physics-based models. In contrast, model training in perfect prognosis downscaling relies on historical data and aims to extrapolate these learned relationships to future conditions. RCM emulation like in EnScale can readily be combined with a state-of-the art bias-correction method to obtain “future observations”. A promising direction for further development is to explore more sophisticated bias correction algorithms that preserve multivariate structure across time, space, and several variables or integrate observational data also in model training [28, 56].
@@
EnScale is computationally efficient and can be re-trained also on other datasets from other GCMs and RCMs or for other regions. Our sparse architecture would allow EnScale to scale to larger domains while still requiring only a reasonable amount of memory. In addition, inference times can be decreased even further with an adjusted data transformation. Also, our emulator could be used to downscale to a higher resolution, for example approximating convection-permitting models as a target. These benefits establish EnScale as a powerful tool for data-driven future climate projections on a local scale and open new opportunities for climate impact studies, where reliable high-resolution projections are essential.
@@ Acknowledgments
Maybritt Schillinger, Maxim Samarin, Nicolai Meinshausen, and Reto Knutti are part of SPEED2ZERO, a Joint Initiative co-financed by the ETH Board. In some instances, ChatGPT was used to rephrase sentences for clarity. We thank Ruth Lorenz and Urs Beyerle for their help with the data-preprocessing and Sebastian Sippel for valuable inputs during the early stages of this project. Further, we benefitted from many helpful and inspiring discussions with numerous people, among them were Antoine Doury, Jose González-Abad, Shahine Bouabid, Sam Allen, Anna Merrifield, Dominik Schumacher, Daniele Nerini, Niklas Boers and his group. We thank Nina Effenberger and Anya Fries for their constructive feedback on the draft which greatly improved the readability of this work. Last but not least, we thank three anonymous reviewers for their careful and constructive feedback on this manuscript.
Open Research Section
All datasets are publicly available from the Earth System Grid Federation (ESGF) infrastructure, via https://esgf-metagrid.cloud.dkrz.de/search. Code is available at https://github.com/m-schillinger/enscale. The repository will be updated to include the final version of the code upon publication.
@@ Appendix A Additional results
A.1 Further samples
@@
Figure A.5: Examples for CorrDiff: Like Fig. 5, but showing examples for the CorrDiff benchmark. These examples are representative for both CorrDiff and CorrDiff-s.
Figure A.6: Examples for EasyUQ: Like Fig. 5, but showing examples for the EasyUQ benchmark.
Figure A.7: Examples for NN-det: Like Fig. 5, but showing examples for the NN-det benchmark.
@@
A.2 Further evaluation of spatial structure
Figure A.11: Power spectral densities
Table A.1: RALSD-avg
Method tas pr sfcWind rsds
NN-det 0.465 8.600 1.390 4.048
EasyUQ 5.959 7.614 8.503 9.969
Analogues 0.150 0.342 0.152 0.169
GAN 1.999 6.894 3.185 4.066
CorrDiff 0.393 0.948 0.432 0.751
CorrDiff-s 0.142 0.526 0.309 0.342
EnScale 0.517 1.299 0.659 1.251
EnScale-t 0.405 1.443 0.573 1.112
A.3 Histograms for benchmark methods
Figure A.12: Histograms for the benchmark methods. As Fig. 7, but showing a selection of benchmarks.
A.4 Further evaluation of extremes
Table A.2: Extreme evaluation regarding marginal quantiles as well as miscalibrationa
Method tas pr
AE $Q^{0.01}_{\text{Wi}}\downarrow$ AE $Q^{0.99}_{\text{Su}}\downarrow$ MCB $Q^{0.1}$ $\downarrow$ MCB $Q^{0.9}$ $\downarrow$ AE $Q^{0.01}_{\text{Wi}}\downarrow$ AE $Q^{0.99}_{\text{Su}}\downarrow$ MCB $Q^{0.1}$ $\downarrow$ MCB $Q^{0.9}$ $\downarrow$
NN-det 2.005 1.366 N/A N/A 0.001 16.963 N/A N/A
EasyUQ 1.094 0.664 0.060 0.064 0.000 6.362 0.028 0.039
Analogues 1.131 0.708 0.103 0.054 0.000 3.852 0.068 0.085
GAN 2.484 1.816 0.097 0.063 0.003 10.481 0.274 0.038
CorrDiff 1.478 1.228 0.056 0.050 0.071 10.345 0.088 0.047
CorrDiff-s 0.825 0.771 0.017 0.019 0.090 3.574 0.030 0.013
EnScale 1.032 0.727 0.018 0.026 0.000 7.972 0.047 0.014
EnScale-t 0.982 1.036 0.017 0.022 0.000 7.472 0.048 0.017
Method sfcWind rsds
AE $Q^{0.01}_{\text{Wi}}\downarrow$ AE $Q^{0.99}_{\text{Su}}\downarrow$ MCB $Q^{0.1}$ $\downarrow$ MCB $Q^{0.9}$ $\downarrow$ AE $Q^{0.01}_{\text{Wi}}\downarrow$ AE $Q^{0.99}_{\text{Su}}\downarrow$ MCB $Q^{0.1}$ $\downarrow$ MCB $Q^{0.9}$ $\downarrow$
NN-det 0.588 1.204 N/A N/A 6.153 12.345 N/A N/A
EasyUQ 0.111 0.333 0.041 0.049 1.155 2.473 0.023 0.038
Analogues 0.106 0.358 0.076 0.095 0.907 2.345 0.064 0.119
GAN 0.266 0.988 0.064 0.068 5.465 9.512 0.091 0.038
CorrDiff 0.264 0.591 0.059 0.043 2.836 7.049 0.051 0.097
CorrDiff-s 0.120 0.649 0.022 0.014 2.138 3.861 0.013 0.033
EnScale 0.136 0.699 0.016 0.012 3.184 10.222 0.015 0.022
EnScale-t 0.167 0.488 0.016 0.016 2.744 10.302 0.014 0.024
A.5 Evaluation of variable dependencies
Firstly, we consider errors in pairwise correlations between variables for three selected variable pairs. For all pairs, the analogue benchmark reaches best or second-best scores. This is because the analogues only reproduce existing RCM data. By construction, the dependencies between variables are correct for the existing data. Scores by EnScale and EnScale-t are slightly higher. Also CorrDiff-s yields low errors, while CorrDiff performs substantially worse. In contrast, EasyUQ which models each variable independently of others, shows much higher errors. This is another example for the benefits of our multivariate approach.
The next metrics evaluates dependencies of the variables conditional on GCM data $X$: we consider calibration of the pairwise difference of variables (last three columns of Tab. A.3, see Sec. 4.2.5 for details). Now, performance of the analogues is worse. Additional evaluation shows that they are again underdispersed (not shown). Instead, EnScale or EnScale-t achieve the best results, followed by CorrDiff-s. This finding again highlights the difference when analyzing marginals vs. conditional distributions, as in Sec. 5.3.5.
Table A.3: Multi-variable evaluationa
Method Correlations Miscalibration
pr-sfcWind sfcWind-rsds tas-pr pr-sfcWind sfcWind-rsds tas-pr
NN-det 0.125 0.062 0.060 N/A N/A N/A
EasyUQ 0.166 0.217 0.055 0.562 0.790 0.388
Analogues 0.026 0.029 0.025 0.355 0.332 0.325
GAN 0.063 0.056 0.042 0.097 0.093 0.123
CorrDiff 0.126 0.094 0.117 0.195 0.194 0.214
CorrDiff-s 0.026 0.046 0.024 0.061 0.052 0.084
EnScale 0.043 0.033 0.035 0.055 0.056 0.069
EnScale-t 0.038 0.027 0.043 0.056 0.057 0.067
A.6 Further time series evaluation
Figure A.13: As Fig. 10, omitting CorrDiff, but presenting all four target variables instead.
Table A.4: Signed and absolute errors in lag-1 autocorrelations (ACF) for each variablea
Method tas pr sfcWind rsds
Error Abs. Error Error Abs. Error Error Abs. Error Error Abs. Error
NN-det 0.01 0.01 0.17 0.17 0.08 0.08 0.11 0.11
EasyUQ -0.03 0.03 -0.07 0.07 -0.15 0.15 -0.03 0.03
Analogues -0.1 0.1 -0.12 0.12 -0.29 0.29 -0.05 0.05
GAN -0.06 0.06 -0.1 0.1 -0.21 0.21 0.02 0.02
CorrDiff 0.01 0.01 0.22 0.22 0.1 0.1 0.09 0.09
CorrDiff-s -0.06 0.06 -0.09 0.09 -0.21 0.21 -0.04 0.04
EnScale -0.08 0.08 -0.09 0.09 -0.22 0.22 -0.05 0.05
EnScale-t -0.01 0.01 -0.01 0.04 -0.02 0.03 -0.01 0.02
a We calculate ACFs in each location for both the samples and the RCM data. We average both the signed difference of samples’ ACF minus target ACF (columns “Error”) as well as the absolute errors in ACFs for samples minus target (columns “Abs. Error.”). Both metrics are averaged across locations.
A.7 Results on extrapolation test period
@@
NN-det-extra 4.4 ± 0.08 240.1 ± 1.83 239.9 0.0 1.51 ± 0.01 1.51 ± 0.01 0.8 ± 0.01
EasyUQ-extra 6.3 ± 0.08 179.6 ± 1.49 305.3 251.3 1.20 ± 0.01 2.14 ± 0.02 6.1 ± 0.02
Analogues-extra 8.0 ± 0.12 249.2 ± 2.91 415.6 332.7 1.74 ± 0.01 1.72 ± 0.01 1.2 ± 0.01
GAN-extra 9.3 ± 0.09 223.3 ± 1.43 377.4 308.1 1.48 ± 0.01 1.45 ± 0.01 2.2 ± 0.01
CorrDiff-extra 6.8 ± 0.11 219.7 ± 2.36 394.1 348.6 1.09 ± 0.01 1.08 ± 0.01 0.9 ± 0.01
CorrDiff-s-extra 7.6 ± 0.13 157.2 ± 1.69 313.9 313.4 1.08 ± 0.01 1.07 ± 0.01 0.9 ± 0.01
EnScale-extra 4.5 ± 0.08 166.8 ± 1.5 343.3 352.7 1.18 ± 0.01 1.19 ± 0.01 1.3 ± 0.01
EnScale-t-extra 4.8 ± 0.08 174.8 ± 1.78 348.2 346.7 1.22 ± 0.01 1.22 ± 0.01 1.2 ± 0.01
@@
NN-det-extra 35.1 ± 0.8 632.3 ± 6.3 632.6 0.0 2.38 ± 0.03 6.41 ± 0.07 10.5 ± 0.13
EasyUQ-extra 54.2 ± 0.93 469.8 ± 5.29 838.8 737.5 1.98 ± 0.02 11.38 ± 0.07 9.6 ± 0.05
Analogues-extra 45.7 ± 0.88 567.2 ± 6.36 935.8 737.9 2.42 ± 0.02 6.30 ± 0.06 5.7 ± 0.07
GAN-extra 53.0 ± 1.24 481.2 ± 5.27 799.1 635.0 1.98 ± 0.02 5.38 ± 0.06 8.8 ± 0.07
CorrDiff-extra 55.1 ± 1.0 438.8 ± 4.93 832.6 788.1 1.90 ± 0.02 4.92 ± 0.05 4.0 ± 0.06
CorrDiff-s-extra 61.7 ± 1.31 434.2 ± 5.11 854.8 842.4 1.85 ± 0.02 4.75 ± 0.05 4.5 ± 0.06
EnScale-extra 36.2 ± 0.81 454.4 ± 5.19 849.1 788.6 1.93 ± 0.02 5.08 ± 0.05 5.4 ± 0.06
EnScale-t-extra 35.5 ± 0.75 455.8 ± 5.38 831.3 750.7 1.92 ± 0.02 5.02 ± 0.05 5.3 ± 0.06
@@
NN-det-extra 2.6 ± 0.03 196.5 ± 1.03 196.6 0.0 1.13 ± 0.01 1.44 ± 0.01 2.0 ± 0.02
EasyUQ-extra 4.2 ± 0.03 141.2 ± 0.78 255.6 229.0 0.88 ± 0.01 2.22 ± 0.01 8.9 ± 0.02
Analogues-extra 3.9 ± 0.04 180.9 ± 1.56 304.1 246.5 1.13 ± 0.01 1.39 ± 0.01 2.0 ± 0.02
GAN-extra 4.7 ± 0.04 151.4 ± 0.84 269.9 236.9 0.95 ± 0.0 1.16 ± 0.01 3.6 ± 0.03
CorrDiff-extra 3.8 ± 0.04 137.3 ± 0.9 269.4 264.2 0.82 ± 0.0 1.00 ± 0.01 1.4 ± 0.01
CorrDiff-s-extra 4.4 ± 0.05 129.2 ± 0.87 255.4 252.4 0.81 ± 0.0 1.00 ± 0.01 1.6 ± 0.02
EnScale-extra 2.6 ± 0.03 133.1 ± 0.84 263.3 260.5 0.85 ± 0.0 1.06 ± 0.01 1.8 ± 0.02
EnScale-t-extra 2.7 ± 0.03 135.2 ± 0.9 264.1 257.8 0.86 ± 0.0 1.07 ± 0.01 1.8 ± 0.01